OpenStudy (idealist10):

Find all solutions of y'=1+x-(1+2x)y+xy^2.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

Looks like a Bernoulli eq to me.

OpenStudy (idealist10):

So how do I work it out?

OpenStudy (anonymous):

\[\begin{align*}y'&=1+x-(1+2x)y+xy^2\\\\ y^{-2}y'&=y^{-2}+xy^{-2}-(1+2x)y^{-1}+x\\\\ -v'&=v^2+xv^2-(1+2x)v+x&\text{setting }v=y^{-1}\\&&\text{then }v'=-y^{-2}y'\\ -v'&=v^2-v+x(v^2-2v+1)\\\\ -v'&=v(v-1)+x(v-1)^2\\\\ -\frac{1}{(v-1)^2}v'&=\frac{v}{v-1}+x \end{align*}\] Try another substitution, like \(t=\dfrac{1}{v-1}\), then \(t'=-\dfrac{1}{(v-1)^2}v'\). \[\begin{align*} t'&=\left(\frac{1}{t}+1\right)t+x\\\\ t'&=1+t+x\\\\ t'-t&=1+x \end{align*}\] which is linear.

OpenStudy (idealist10):

Thank you!